OpenStudy (anonymous):

Harry rides his bike 25 & 3/8 miles. Karen rides 25 & 2/5 miles. Who rides farther? Please help! WILL GIVE MEDAL!

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

hi! so the 25 mile is the same for both riders right? so what we need to compare here?

OpenStudy (texaschic101):

They both have the 25 miles....so compare the fractions... 3/8 = 15/40 2/5 = 16/40 does that help ? Can you figure it out now ?

OpenStudy (anonymous):

@texaschic101 can you help a little more im so confused

OpenStudy (anonymous):

Omg im so dumb, yes that helps so much thank you!!